OpenStudy (anonymous):

do anyone know the distance formula for (-5,2) and (-9,9)

OpenStudy (akashdeepdeb):

Distance formula is not specific to 2 points. Understand this. YOU NEED TO FIND THE **DISTANCE** between the 2 points. Distance formula is like this \[\sqrt{(x2-x1)^2 + (y2-y1)^2}\] For any 2 points (x2,y2) = [here] = (-5,2) (x1,y1) = [here] = (-9,9) Put them in the distance formula and you'll get the distance. :) Understood? :)

OpenStudy (anonymous):

yes working it out now

OpenStudy (anonymous):

@AkashdeepDeb i got -65

OpenStudy (akashdeepdeb):

I think you got \[\sqrt 65\] And that is the distance.

OpenStudy (anonymous):

18 13 8 11 these are the choices

OpenStudy (akashdeepdeb):

It is APPROXIMATELY 8. \[\sqrt 65 \approx 8 = \sqrt 64 \] Understood? :)

OpenStudy (anonymous):

oo ok thnkss

OpenStudy (akashdeepdeb):

:)

OpenStudy (anonymous):



OpenStudy (anonymous):

@AkashdeepDeb

OpenStudy (akashdeepdeb):

Use trigonometry.

OpenStudy (akashdeepdeb):

Or the special triangle ratios.

OpenStudy (anonymous):

tan30/4??